OpenStudy (anonymous):

A solution that is 80% water is mixed with another solution that is 90% water. The resulting solution is 88% water. What is the ratio of 80% solution to 90% solution in any volume of the 88% solution? a.1:5 b. 1:4 c. 4:5 d. 8:9

OpenStudy (shinalcantara):

let 'x' be the amount of solution with 80% water 'y' be the amount of solution with 90% water you'll have the equation as: 0.80x + 0.90y = 0.88(x+y) 0.80x + 0.90y = 0.88x + 0.88y 0.90y - 0.88y = 0.88x - 0.80x 0.02y = 0.08x divide the whole equation with 0.02 1y = 4x therefore you'll have the ratio x:y = 1:4

OpenStudy (shinalcantara):

*x:y = 4:1